UT 2019 Dec 03 - Eng Half Night

Plan:

  • LUCI-2/AGW-4 confirmatory transform - highest priority - requires seeing <1.2" on the DIMM
    • DONE AGW4 - Test IDL config parameters for IDL off-axis collimation - 0 time required (parallel to other LUCI2 activities) [John]
    • DONE (done, but perhaps not resolved) Rotator ZP: Scripts for INAF and LBTB masks (including new LMS versions and fraternal LBTB scripts) located in:
      • /home/observer/share/Service_Observing/Eng_Dec_2019/LUCI_RT_ZP on RM507-5/4 computer
      • /home/observer/Service_Observing/Eng_Dec_2019/LUCI_RT_ZP on RM507-1-3computer
  • Photometric Standard stars (any seeing - only if clear) - LUCI, LBC, MODS (priority order)
    • LUCI: Zeropoints (clear - any seeing) - low priority - requires min 2 airmass visits - all night
      • /home/observer/share/Service_Observing/Eng_Nov_2019/LUCI_ZP/ RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/LUCI_ZP/ RM507-1-3
  • LBCB+LUCI2: Verify mixed-mode use with dithering and guiding (all combos)
    • can be done in any conditions
  • DONE PCS patch for IT 6940
    • Activate patch (cd /lbt/tcs/current && ./tcs-build latest PCS)
    • send a variety of presets including NS (doesn't need to be targeted for the patch, just use PCS as usuall)
    • (sky optional)
  • LBC
    • DONE Superfoc sequences (at least in the default filters) at 0, 180 deg position angle - 1 hour, priority medium - requires good seeing but not clear skies
      • lbto@obsN:~/Startup/Startup_Sep_2019/LBC
    • DONE PtAutoadjust software test
  • Poor Weather Backup
    • H-band Library- Bino LUCI - from handover - good through morning twilight
      • /home/observer/share/Service_Observing/Eng_Nov_2019/H-band_Library on RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/H-band_Library on RM507-1-3
    • Arp 187 LS - Bino LUCI - observable (start collecting data) until 2:15am (local time) -
      • /home/observer/share/Service_Observing/Eng_Nov_2019/LocalMergers on RM507-5/4
      • /home/observer/Service_Observing/Eng_Nov_2019/LocalMergers on RM507-1-3
    • DONE (one target observed) Local Mergers CO-Bandhead - Bino-LUCI - targets span all 2nd half of the night
      • Any conditions - 3 dither positions if bino, 6 if monocular -longer if really thick clouds or poor seeing
      • README has priority list
      • /home/observer/share/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - on RM 507-5/4
      • /home/observer/Service_Observing/Eng_Dec_2019/Poor_Weather_Backup/Local_Mergers - on RM 507-1-3

Log

OSA: S. Allanson
Observers: J. Hill
O. Kuhn

Summary

We did not verify the LUCI2/AGw4 transform, but obtained some useful data.
  • We aligned two masks to confirm the rotator zeropoints measured for both L1/AGW1 and L2/AGw4 on Nov 18 UT. In particular, both of these masks suggested a 0.45 (or perhaps more accurately, a 0.47 deg) offset needed for L1, similar to what was found in Nov. This is pretty large and would affect programs that aim to align the slit along the PA of a galaxy or between two targets, or a target + a brighter reference star. Decreasing the LFBG rotator zeropoint by 0.45 deg resulted in a smaller dtheta - this LFBG rotator zeropoint was not left in place, but can be referred to if there is an opportunity to update the transform for L1.
  • We obtained long-slit LUCI1+LUCI2 spectra for a D target
  • With the LBCs, we confirmed that the ptautoadjust now installed on obsN works for both SX and DX.
  • We obtained superfoc sequences with both LBCs in the reference filters and also squeezed in a sequence at Us and R-Bessell
  • We obtained twilight images (not flats - clouds) at i-SLOAN which showed the effect of the spots on the filter, which were noted by Elliott and were there in the Oct 22 UT flats as well. These should be cleaned if possible.
  • Activated PCS patch 20191202 that should fix IT 6940, however because the probability of running into the error was very low already before the patch, we did not try to find failures. We did, however, send many binocular LUCI and LBC presets without problems.
  • John fixed a problem with LBTplot.py (python2 version) not updating the plot under Fedora. John also started working on the python3 version LBTplot_new.py.

Details

parallel tests of IDL active optics for PEPSIPFU during first half

running IDL active optics in parallel with GCS using iif_register, /RPFU and gcs_wfsc_display. Initially, RPFU in IDL had the wrong sign for Z7,Z8,Z9,Z10. LPFU signs were OK.

Use print, station to show the currently selected focal stations in IDL. Say retall after a ^C to get back to main program variables. Use print, station_params to see the list of possible stations. As we found on 20191218, station_check.pro must not find more than one instance of agwname in the list of stations (from station_define.pro).

02:45 UT with RPFU: adding 180 deg to wfsc.rot_offset fixed the signs of Z7-Z10. The scaling and rotation of Z9, Z10 are still a bit off from GCS.
IDL for RPFU and LPFU is still computing Z12-Z21.
IDL seems to want more negative Z11 than GCS for RPFU.
This trick of changing wfsc.rot_offset only works on-axis. See log below where is messes up the off-axis field corrections.

In LBTplot.py (python2 version on obs1), I replaced gobject.timeout_add_seconds with glib.timeout_add_seconds to get plot refreshing working under Fedora. I also fixed a ascii template problem in wplot_newt.pro related to changing tcslog information. I commented out FITEXY for guideplot_hdf.pro because of Fedora IDL path issues for the astron package.

07:05 Handover from PEPSIPFU observing - Cirrus, but not too bad. Seeing was very variable during the first half, from ~1-1.6", but close to handover it might have been more stable at 1".

LUCI Rotator Zeropoints

07:09 Reconfiguring and moving masks into L1 and L2. Will check rotator zeropoints. Currently, they are:
L1 LFBG RZP -187.315
L2 RFBG RZP 176.85

07:10 Steve activated PCS Patch and it worked OK for the subsequent presets.

07:14 Point and Collimate at field of KVVDSFin LBTB Mask 1 (want to get LMS files that have added stars, in ~/Service_Observing/Eng_Dec_2019/LUCI_RT_ZP):
name of field L1 mask ID LMS file L2 mask ID LMS file comment
LBTB Mask 1 901807.HVVDSFin luci1.7.HVVDSFin.lms 994220.KVVDSFin luci1.6.KVVDSFin.lms added stars (6 alignment stars)
LBTB Mask 1 901807.HVVDSFin luci1.93.HVVDSFin.lms 994220.KVVDSFin luci1.89.KVVDSFin.lms original (4 alignment stars)
LBTB Mask 2 905739 luci1.8.HVVDSFin.lms 905739 luci1.4.KVVDSFin.lms added stars (6 alignment stars)
LBTB Mask 2 905244 luci1.97.HVVDSFin.lms 905739 luci1.98.KVVDSFin.lms original (5 alignment stars)

07:29 Right side needed some help to converge. Finally done and slewing to LBTB Mask 1 field.
  • L2 Camera error. No limit hit. Moved to N1.8, then init and then back to N3.75. OK now.
  • Meanwhile, collimation on right side falling apart. We are low (ZD 52), but that should not be a problem... John and Steve are investigating while I take data on the left side.
  • John and Steve think that the problem is with the flat (not very many modes used), but they will need to go to a higher elevation to reset it.

07:37 L1 acq images:
  • 0002 (sky) but with blinds, because first obs item did not complete.
  • 0003 (obj) same - blinds
  • 0004 (mask) number is off the top, but ref slits clearly visible and there's a slit very close to bottom edge.
  • 0005 (sky again at -10, -10)
  • 0006 (obj at 0,0)
  • Alignment: I can see where stars should be but the offsets it is calculating and the errors are enormous, ~18". If I remove stars 34 and 35, then it is better:
    • Using stars 30-33, xshift = 11.7, yshift = 9.25, angle is 0.42769. X-error is green, Y-error is yellow.
    • using stars 30, 31 and 33, xshift = 11.60, yshift = 9.37 and angle = 0.431. Errors are green. Adopting this result and sending.
    • Error: could not insert item into scheduler. I ran RTD separately using /lbt/lucifer/lcsp/bin/startRTD.sh 1
    • Ran RTD from observer panel now and redid the mos alignment. Had to manually tell it where stars 30-33 were. With these 4 selected, the errors on 34 and 35 were ~8". Excluded 34,35 from the fit. Results with stars 30-33 are: xshift = 11.4, yshift = 9.29, angle =0.419. Errors are green (0.07, 0.09 arcsec)
  • 0007 is junk (confirmatory image, but mask was in turnout)
  • 0008 is with mask in FPU, but I see nothing in the boxes, even with a closer look in ds9.

Mask alignment summary:
  • RTD launched from xterm on 32-bit machine (obs2) does not insert offset item to sequencer; need RTD launched from observer panel.
  • RTD launched from xterm and from observer panel yield different answers for alignment.
  • We want to move to 64-bit anyway, but meanwhile we need to check this more for observers at the summit - is /lbt/lucifer/lcsp/bin/startRTD.sh 1 not the correct command or path?

08:12 Slew to higher elevation to set flat. That gives more modes. Active Optics converged near zenith and also at ZD 48 deg and off-axis. Now trying a different probe position. At high elev, it converges well, but at low elevation, convergence is not optimal.

08:36 Going to UGC5101 --- the transform is not quite ready yet and the field for a second mask alignment has not yet risen.
  • bino luci telluric
    • field images are: L1 0009 and L2 0003
    • 0.5 slit: L1 0010 and L2 0004
    • sending offsets and taking confirmatory images: L1 0011 and L2 0005
      • 08:46 collimation is starting to fall apart on DX side. Guide star FWHM now 2.5"
    • taking spectra: L1 0012-00 and L2 0006-
      • a big diff in central where central wavelength falls for L1 and L2, about 300 pix.
  • bino ugc5101
    • field images are: L1 15, 16 and L2 09, 10
    • slit images are: L1 17 and L2 11
    • offsets to align target on slit and confirm through slit images are: L1 18 and L2 12
    • spectra started at 09:11, L1 19-21 and L2 13-15
      • seeing is good, 0.7-0.8" on guiders.
      • on images L1 20 and L2 14, the guide probe is covering the position of the spectrum on images 19 and 13.
    • launching a version with smaller dithers: this will start with images L1 21 and L2 15
    • L1 21-24 and L2 15-17
      • Tried to subtract 19 from 22 in aladin - WCS must be intefering with the subtraction - crazy. RTD launched from Observer Panel on obs2. On rm507-4, better result, but then on a LUCI2 pair (16 and 17), on rm507-4, WCS is again interfering. There is a vertical shift before subtraction.
10:30 A 2nd MOS alignment (INAF mask 3):
INAF mask 3 924235 luci1.2.maskz2.lms same for L1 & L2, 4 alignment stars
Alignment images (sky, object, mask) are: L1 (24,25,26) and L2 (18,19,20) Results are:
  • L1 x-shift = 1.35, y-shift = -0.79, dtheta = 0.466 with 4 stars. Errors in x 0.088 (green) and y 0.119 (yellow). With 3 stars (30,32,33): x-shift=1.24, y-shift=-0.85, dtheta=0.42 deg and errors both green x 0.03, y 0.04 arcsec.
  • L2 x-shift = -0.87, y-shift=0.068, dtheta=0.0035 deg with 4 stars. Errors in x 0.072 and y 0.067 arcsec.
  • Confirm thru-mask images L1 27 and L2 21 - stars are in boxes.

L1 RZP does seem to be off by ~0.45 deg. Current RZP = -187.315; Try RZP - dtheta = -187.765 deg.

10:58 Steve restarted PCS after I did the following.
[tcs@tcs1 PCS]$ cp PCSInstrument.conf PCSInstrument.conf.20191203
[tcs@tcs1 PCS]$ vi PCSInstrument.conf 
[tcs@tcs1 PCS]$ pwd [tcs@tcs1 PCS]$ diff PCSInstrument.conf PCSInstrument.conf.20191203
203c203
< LEFTZEROPOINT      -187.765
---
> LEFTZEROPOINT      -187.315

11:01 Reloaded script and sent preset to same field, INAF mask 3.

Alignment images are: L1 (28,29,30) and L2 (22,23,24).

Results are:
  • L1 x-shift = 3.33, y-shift=-1.31 dtheta=0.01778 deg with 4 stars. Errors are x 0.098, y 0.116. With 3 stars (30,32,33): x-shift=3.36, y-shift=-1.26, dtheta=0.0136 deg. Errors are x 0.077, y 0.064 arcsec.
  • L2 x-shift = -0.93, y-shift=0.178, dtheta=-0.00085 with 4 stars. Errors are x 0.0666 y 0.0616
  • The obsitem with the L1 offsets went into the normal place, but the obsitem with the L2 offsets went above the preset ??? Weird.

Repeat: New alignment images are: L1 (31,32,33) and L2 (25,26,27).
  • L1 x-shift = 3.37, y-shift=-1.35 dtheta=0.0166 with 4 stars. 3.4, -1.33, dtheta=0.013 with 3 stars and errors 0.075, 0.06.
  • L2 x-shift = -0.88, y-shift=0.15, dtheta=0.0058 with 4 stars. Errors 0.065, 0.054.
  • Confirmatory images: L1 34 and L2 28.

11:30: Reverted old PCSInstrument.conf and reconfigured to LBCB+LBCR.
[tcs@tcs1 PCS]$ cp PCSInstrument.conf PCSInstrument.conf.20191203_correct_LUCI1ZP
[tcs@tcs1 PCS]$ cp PCSInstrument.conf.20191203 PCSInstrument.conf
[tcs@tcs1 PCS]$ diff PCSInstrument.conf.20191203_correct_LUCI1ZP PCSInstrument.conf
203c203
< LEFTZEROPOINT      -187.765
---
> LEFTZEROPOINT      -187.315

parallel tests of IDL active optics for AGW4

John did this work in parallel with Olga's rotator zeropoint measurements.

07:26 The RFBG wfs pupil cutout is one row too low for the off-axis guide star at x=0, y=532.4.

07:43 Try changing pupil_shift_y from +1 to -1. That was obviously wrong (made things worse).

07:45 Try changing pupil_shift_x from -1 to +1. That was obviously wrong (made things worse).

07:49 Change WFS_hotspot_y from 236.8 to 223.5 in both GCSR and IDL. That's the value that works for this off-axis star.

07:55 Changed wfsc.rot_offset from 0 to 180 deg for RPFU (while undoing the forced signs for Z7-Z10). This is successful in getting rid of the spurious signs we added last month, but now there is a disagreement about the field corrections. "Rotated Zernikes" are the same for either configuration. This can be a hint for RPOL where we were having issues with the Field Corrections.

08:05 revert to the previous IDL configuration in agw_define.pro (forced signs, and rot_offset=0).

08:12 This different off-axis star wants the original pupil position. Do we have a pupil wobble issue?

08:16 On-axis collimation wants us to go back to the original pupil_center_y of 236.8.

08:30 Off-axis star at x=0, y=462. The RFBG computed pupil shift seems off again. IDL and GCS agree on the shift (see flips earlier).

08:37 Off-axis star at x=-56, y=597 seems like the pupil cutout is wrong by one lenslet row again.

08:41 Change pupil shift scale factor from 0.19 to 0.17. That didn't seems to change anything, because it was only 1.5 pixels.

08:46 Change pupil shift scale factor from 0.19 to -0.19. That was clearly wrong, so back to +0.19.

09:00 Off-axis star at x=27 y-535. Put WFS_hotospot_pupil_y back to 223.5, because that is better for the off-axis stars, and we don't do science with on-axis guide stars.

09:40 verified that Field Corrections are the same on both sides.

LBC tests

11:55 Sending a bino LBC preset. Failed at first with error messages about bad polys (poly posn beyond max limit). Second try was successful.

11:58 Running dohybrid. Glass is about 0.4-0.6 deg C below ambient temperature (Glass T~6 C and Ambient T~ 6.6 for sx and 6.5 for dx). Numbers below show that WRS corrected mostly astigmatism, and then FPIA corrected both astigmatism, essentially backing out what WRS introduced, and coma. WRS:
(Z4) (Z5) (Z6) (Z7) (Z8) (Z11) (Z22) seeing WIQ  
-6440 630 1834 -74 70 261 0 1.48 2.61 red
-4132 -464 -4270 21 -131 98 0 0.52 2.06 blue

FPIA, where the 1st 7 columns are the same as above:
  -875      0      0      0      0    500      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  3483   -268   -792  -2235    331   -258    -89    y  0.68  1.73                      /newdata/lbcr.20191203.120305.fits  r-SLOAN 
  -813   -274   -272   -116   -153     87      0    y  0.60  0.36                      /newdata/lbcr.20191203.120447.fits  r-SLOAN 
    55    -73   -223    -21    -72     12    -13    y  0.65  0.11                      /newdata/lbcr.20191203.120616.fits  r-SLOAN 
    72   -208      5     75     20     61      1    y  0.71  0.12                      /newdata/lbcr.20191203.120743.fits  r-SLOAN 
   565      0      0      0      0   -323      0    y  0.00  0.00                     dofpia_sph3_adjustment             
 -1050      0      0      0      0    600      0    y  0.00  0.00                     dofpia_sph3_adjustment             
  1624   -344   1025    -16  -1007    -94     -5    y  0.91  0.85                      /newdata/lbcb.20191203.120309.fits  V-BESSEL
   319   -271    650     72   -194   -140      0    y  0.93  0.29                      /newdata/lbcb.20191203.120451.fits  V-BESSEL
   -26   -101    455     33      9    -90      0    y  0.99  0.17                      /newdata/lbcb.20191203.120621.fits  V-BESSEL
  -161    -62    285    -31    106     49     12    y  1.05  0.15                      /newdata/lbcb.20191203.120746.fits  V-BESSEL
   820      0      0      0      0   -469      0    y  0.00  0.00                     dofpia_sph3_adjustment         

12:01 Ran CP script and ran ptautoadjust on each of the images. Before doing so, set appropriate instrument in LBTtools. ptautoadjust correctly worked on each side.
  • CP script was not pointing at a star, but it was OK for this test.

12:20 Ran CP script, now edited, to do actual co-pointing and range-balancing (lbcrangebal).

12:26 Collimate with FPIA prior to starting superfoc sequences. DIMM seeing is sub-arcsec.

12:30 Superfoc at PA=0 using reference filters V and r.
lbto@obs2:12 % gethead /newdata/lbcr.20191203.12*.fits -x 2 filter focusoff pa_pnt
lbcr.20191203.123041.fits[2] r-SLOAN -0.144 -0.02992
lbcr.20191203.123124.fits[2] r-SLOAN -0.084 -0.02344
lbcr.20191203.123206.fits[2] r-SLOAN -0.024 -0.02324
lbcr.20191203.123249.fits[2] r-SLOAN 0.036 -0.02476
lbcr.20191203.123330.fits[2] r-SLOAN 0.096 -0.01875
lbcr.20191203.123413.fits[2] r-SLOAN 0.156 -0.02155
lbcr.20191203.123453.fits[2] r-SLOAN 0.216 -0.02208
lbto@obs2:13 % gethead /newdata/lbcb.20191203.12*.fits -x 2 filter focusoff pa_pnt
lbcb.20191203.123045.fits[2] V-BESSEL -0.200 -0.02893
lbcb.20191203.123133.fits[2] V-BESSEL -0.140 -0.02246
lbcb.20191203.123222.fits[2] V-BESSEL -0.080 -0.02145
lbcb.20191203.123312.fits[2] V-BESSEL -0.020 -0.02161
lbcb.20191203.123403.fits[2] V-BESSEL 0.040 -0.02748
lbcb.20191203.123454.fits[2] V-BESSEL 0.100 -0.02105
lbcb.20191203.123542.fits[2] V-BESSEL 0.160 -0.02123
Followed immediately by series at PA=180:
lbcr.20191203.123808.fits[2] r-SLOAN -0.144 179.94707
lbcr.20191203.123850.fits[2] r-SLOAN -0.084 179.98123
lbcr.20191203.123933.fits[2] r-SLOAN -0.024 179.97975
lbcr.20191203.124020.fits[2] r-SLOAN 0.036 179.98042
lbcr.20191203.124102.fits[2] r-SLOAN 0.096 179.98417
lbcr.20191203.124145.fits[2] r-SLOAN 0.156 179.98331
lbcr.20191203.124227.fits[2] r-SLOAN 0.216 179.98176
lbcb.20191203.123810.fits[2] V-BESSEL -0.200 179.94821
lbcb.20191203.123859.fits[2] V-BESSEL -0.140 179.97918
lbcb.20191203.123948.fits[2] V-BESSEL -0.080 179.98515
lbcb.20191203.124037.fits[2] V-BESSEL -0.020 179.97727
lbcb.20191203.124126.fits[2] V-BESSEL 0.040 179.97734
lbcb.20191203.124214.fits[2] V-BESSEL 0.100 179.98049
lbcb.20191203.124304.fits[2] V-BESSEL 0.160 179.97884
and one more at PA=0.

12:53 Now a series using Us and R-Bessel filters (12:53-12:58)

12:58 Followed by another at Vr.

13:05 Well into twilight but will try a new field a higher up.

13:07 Started FPIA before slew had finished and then cntrl-C'd and sent backout twice when it wasn't even needed once.

13:11 Started rV PA=0 superfoc (skipped lbcrangebal). Too light. But also too cirrusy for sky flats. (Would have liked i-SLOAN).

13:18 taking some quick snaps at g-SLOAN and i-SLOAN to evaluate effect of spots on i-SLOAN. I think the spots are still on i-SLOAN - there is a circle between chips 2 and 3. See, e.g. lbcr.20191203.132209.fits.

13:26 getting very cloudy, but, while we are configured for LBCs and looking at sky, trying to run through closed dome script that takes an image through each of the filters.

14:00 Closed up. Took some LBC biases (LBCR biases have horizontal stripes) and LUCI Cals.

-- BarryRothberg - 26 Nov 2019
Topic revision: r11 - 04 Dec 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback